A compound contains 174.86 grams of iron and 74.14 grams of oxygen. Construct the empirical formula.
1
Expert's answer
2020-01-13T03:33:00-0500

Ar(Fe)=55.85


n(Fe)=m/Ar=174.86/55.85=3.13


Ar(O)=16.0


n(O)=74.14/16=4.63


n(O)/n(Fe)=4.63/3.13=3/2, the formula is Fe2O3